OpenStudy (anonymous):

Multiply. 2/24 multiplied by 16.

OpenStudy (diyadiya):

4/3

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

\[\frac{2}{24}*16=2*\frac{2}{3}=\frac{4}{3} \]